Business and finance. When x number of units are sold, the price of each unit (in dollars) is given by:
p= -x/2 + 75
Find the unit price when the following quantities are sold: 2, 7, 9, 11.

Answer by rapaljer(4671)

p= -x/2 + 75

If x=2,
p=-2/2+75
p=-1+75=74

If x=7,
p=-7/2+75
p=-3.5+75=71.5

If x=9,
p=-9/2+75
p=-4.5+75=70.5

If x=11,
p=-11/2+75
p=-5.5+75=69.5

R^2 Retired from SCC